OpenStudy (anonymous):

x^2 - 6x - 7 in vertex form

OpenStudy (anonymous):

complete the square

jigglypuff314 (jigglypuff314):

Hello @attababe33 and Welcome to OpenStudy! :) to make it into vertex form, you need to complete the square that means making the equation look something like x^2 - 6x + 9 to get it equal to something like (x - 3)^2 so what would you add to your given equation to make it look like x^2 - 6x + 9 ? :)

OpenStudy (anonymous):

2

jigglypuff314 (jigglypuff314):

mmm not quite @attababe33 because you have a -7 there what would you have to add to -7 to get +9 ? :)

OpenStudy (anonymous):

oh 16 ? o:

jigglypuff314 (jigglypuff314):

yes! :) good, but just adding 16 would unbalance things so we have to subtract a 16 too x^2 - 6x - 7 + 16 - 16 = x^2 - 6x + 9 - 16 = (x^2 - 6x + 9) - 16 now turn the (x^2 - 6x + 9) into a (x - 3)^2 what would you get overall? :)

jigglypuff314 (jigglypuff314):

@attababe33 ? :)

OpenStudy (anonymous):

I have no idea :( its hard to understand

jigglypuff314 (jigglypuff314):

sorry, I'll try to explain it better what part do you not understand?

OpenStudy (anonymous):

are you asking me to put it like (x-3)^2 - 16?

jigglypuff314 (jigglypuff314):

yes! exactly! :D

OpenStudy (anonymous):

what would i have to do to find the answer?

jigglypuff314 (jigglypuff314):

well, that would depend on the vertex form that your teacher wants there is y = (x-h)^2 + k or y - k = (x-h)^2 both are considered "vertex form"

OpenStudy (anonymous):

y= (x-h)^2 + k (: do I just plug it in?

jigglypuff314 (jigglypuff314):

well you've already got "(x-3)^2 - 16" ;) y = (x-3)^2 - 16 would be your answer

OpenStudy (anonymous):

I feel really dumb. I forgot how to do all of it, thank you so much! :D

jigglypuff314 (jigglypuff314):

hehe I'm glad I could help :D